x2+12x-3 what value should be on each side to complete the square?
Answer by stanbon(75887)

x2+12x-3 what value should be on each side to complete the square?
----
x^2+12x-3 = y
----
x^2+12x + ? = y + 3 + ?
---
Complete the square:
x^2 + 12x + 6^2 = y + 3 + 6^2
----
Factor:
(x+6)^2 = y + 39
etc.
----
Answer to your question: 6^2 or 36
===================================
Cheers,
Stan H.